# Luminis Stellarum et Machina: Applications of Machine Learning in Light Curve Analysis

### v1

@@ Abstract
The rapid advancement of observational capabilities in astronomy has led to an exponential growth in the volume of light curve (LC) data, presenting both opportunities and challenges for time-domain astronomy. Traditional analytical methods often struggle to fully extract the scientific value of these vast datasets, especially as their complexity increases. Machine learning (ML) algorithms have become indispensable tools for analyzing light curves, offering the ability to classify, predict, discover patterns, and detect anomalies. Despite the growing adoption of ML techniques, challenges remain in LC classification, including class imbalance, noisy data, and interpretability of models. These challenges emphasize the importance of conducting a systematic review of ML algorithms specifically tailored for LC analysis. This survey provides a comprehensive overview of the latest ML techniques, summarizing their principles and applications in key astronomical tasks such as exoplanet detection, variable star classification, and supernova identification. It also discusses strategies to address the existing challenges and advance LC analysis in the near future. As astronomical datasets continue to grow, the integration of ML and deep learning (DL) techniques will be essential for unlocking the full scientific potential of LC data in the era of astronomical big data.
@@ 1 Introduction
Time domain astronomy (TDA) is a rapidly evolving field, driven by the continual discovery of new phenomena and the expansion of observational capabilities Ball & Brunner (2010); Pesenson et al. (2010). This field investigates time-varying characteristics of celestial objects using data from diverse messengers, including gravitational waves, neutrinos, and electromagnetic radiation across various photon-energy bands Vaughan (2013). With the advent of advanced wide-field, multi-epoch sky surveys such as the Sloan Digital Sky Survey (SDSS) York et al. (2000), Zwicky Transient Facility (ZTF) Graham et al. (2019), Panoramic Survey Telescope and Rapid Response System (Pan-STARRS) Kaiser et al. (2010), and the upcoming Vera C. Rubin Observatory (formerly LSST) Ivezić et al. (2019), the volume of transient discoveries has grown exponentially. For instance, ZTF generates alerts for over 100,000 events each night, while LSST is projected to surpass this by an order of magnitude. These alerts, which indicate significant flux density changes or new spatial positions, represent invaluable data streams for studying transient phenomena such as supernovae, pulsars, and gamma-ray bursts. TDA focuses on analyzing time-series data to construct light curves (LCs), which depict the variation in brightness of celestial objects over time and are fundamental to understanding the physical processes driving these phenomena.
@@
Machine learning (ML) has become an essential tool for automating data classification, archiving, and retrieval Fluke & Jacobs (2020); Baron (2019). Unlike traditional feature extraction methods, which require domain expertise, deep learning (DL) automatically learns data representations. As a subset of ML, DL has gained widespread popularity and has been successfully applied in various domains, including speech recognition, image processing, and natural language processing Nassif et al. (2019); Deng & Li (2013); Rodellar et al. (2018); Mishra & Celebi (2016); Ussipov et al. (2024a, b); Zhao et al. (2019); Pathak et al. (2018); Ezugwu et al. (2022); Ahuja et al. (2020); Otter et al. (2020); Young et al. (2018). In astronomy, ML has emerged as a powerful tool for automating the classification and analysis of LCs, enabling researchers to extract insights from vast datasets. Despite the growing adoption of ML techniques, there remains a need for a systematic review of algorithms and models specifically tailored for LC classification, particularly in addressing challenges such as class imbalance, noise, and interpretability. Interpretability is crucial in astronomical applications, as it ensures that ML models not only provide accurate predictions but also offer insights into the underlying physical processes, fostering trust and enabling scientific discovery.
@@
This article provides a comprehensive overview of ML algorithms and DL models in LC analysis, covering applications such as exoplanet detection, variable star classification, and supernova identification. By presenting a structured and insightful examination of ML techniques for LC analysis, this review contributes to a deeper understanding of current methodologies and offers guidance for their effective application in future astronomical research.
@@
The rest of the paper is organized as follows. Section 2 provides an overview of major photometric data surveys relevant to LC analysis. Section 3 introduces fundamental ML concepts and architectures. Section 4 explores key applications of ML techniques to LC analysis. Section 5 outlines current challenges and open research questions in the field. Finally, Section 6 concludes with a summary and outlook.
@@ 2 Photometric Survey Datasets
Over the past few decades, large-scale photometric surveys have transformed time-domain astronomy, providing vast, multi-wavelength datasets that are readily accessible through online databases. These surveys have been essential in detecting and classifying variable stars, discovering exoplanets, and identifying transient astronomical events. As shown in Figure 1, the data volume from astronomical surveys has grown exponentially, approximately doubling every 16 months. The combination of different survey datasets enables comprehensive studies of stellar variability across various timescales and wavelengths, driving both traditional astrophysical research and modern ML applications.
@@
Photometric surveys can be broadly categorized into ground-based and space-based missions. Ground-based surveys provide wide-field and high-cadence monitoring, often covering extensive regions of the sky over long time periods. In contrast, space-based missions offer high-precision, uninterrupted observations unaffected by atmospheric distortions, enabling the detection of minute variations in stellar brightness.
Table LABEL:table1 provides an overview of major ground-based and space-based photometric surveys, including their active years, observed bands, sky coverage, and references.
2.1 Ground-based Observatories
Ground-based photometric surveys have significantly expanded understanding of stellar variability, exoplanetary systems, and the dynamic nature of the night sky. Through wide-field imaging and long-term monitoring, these surveys detect brightness variations in stars, uncovering key processes involved in stellar evolution. Covering different sky regions and wavelength bands, they provide extensive datasets that support various astrophysical studies. Below, ground-based observatories are grouped based on their primary target regions.
  • All-sky: ASAS, ROTSE, WASP, WISE, ASAS-SN
  • Milky Way (MW): MACHO, OGLE, UKIDSS
  • Wide field: LINEAR, CRTS, HATnet, VVV, HiTS, NGTS, Pan-STARRS, WFST, Rubin Obs.
  • Northern sky: NSVS, ZTF
  • Stripe 82: SDSS
2.2 Space-based Missions
Space-based photometric surveys have revolutionized precision astronomy by eliminating atmospheric distortions and enabling uninterrupted observations. These missions provide stable, high-sensitivity measurements, allowing for the detection of minute brightness variations in stars. Their contributions extend beyond exoplanet discovery to studies of stellar oscillations, galaxy evolution, and cosmic distance measurements. The extensive datasets from these surveys not only refine astrophysical models but also support ML applications in time-series analysis, enabling automated classification of variable stars, exoplanets, and other celestial phenomena. Below, space-based observatories are grouped based on their primary target regions.
  • All sky: Hipparcos, Gaia, TESS
  • Milky Way (MW): CoRoT, Kepler
  • Wide field: PLATO, ULTRASAT, Roman ST
  • Targeted field: JWST
@@ 3 Machine learning fundamentals
Figure 2: Taxonomy of machine learning approaches.
Figure 3: Basic process of machine learning.
ML enables systems to autonomously learn patterns from data and improve decision-making through experience, mirroring aspects of human cognition. Unlike traditional astronomical programming, which relies on explicit physical rules, ML algorithms derive implicit relationships directly from observational data. This data-driven approach offers flexibility for solving complex, nonlinear problems that defy conventional analytical methods Rodríguez et al. (2022); Kembhavi & Pattnaik (2022); Sen et al. (2022).
ML algorithms can be broadly categorized into supervised and unsupervised methods, often referred to as predictive and descriptive, respectively. These approaches may also be combined to form semi-supervised methods. Figure 2 provides a visual taxonomy of these approaches. Supervised algorithms learn mappings between input features and predefined target variables using labeled training data curated by domain experts (see e.g., Connolly et al. (1995); Collister & Lahav (2004); Reis et al. (2018a); Daniel et al. (2011); Fiorentin et al. (2007); Richards et al. (2012); Laurino et al. (2011); Masci et al. (2014); Morales-Luis et al. (2011); Bloom et al. (2012); Djorgovski et al. (2016); Mahabal et al. (2008); Miller (2015); Brescia et al. (2012); Krone-Martins et al. (2014); Wright et al. (2015); Lochner et al. (2016); D’Isanto et al. (2016); Castro et al. (2017); Naul et al. (2018); Ishida et al. (2019); D’Isanto & Polsterer (2018); Zucker & Giryes (2018); Delli Veneri et al. (2019); Krone-Martins et al. (2018); Mahabal et al. (2019); D’Isanto et al. (2018); Norris et al. (2019)). Unsupervised methods, conversely, autonomously identify hidden structures or relationships within unlabeled datasets. These techniques are typically divided into three subcategories: clustering (grouping similar data points), dimensionality reduction (extracting salient features), and anomaly detection (identifying outliers) (e.g., Boroson & Green (1992); D’Abrusco et al. (2009); Protopapas et al. (2006); Vanderplas & Connolly (2009); Ascasibar & Sánchez Almeida (2011); Almeida et al. (2010); D’Abrusco et al. (2012); Meusinger et al. (2012); Krone-Martins & Moitinho (2014); Fustes et al. (2013); Baron et al. (2015); Hocking et al. (2015); Nun et al. (2016); Gianniotis et al. (2016); Baron & Poznanski (2017); Reis et al. (2018b, c)). Anomaly detection holds particular promise for astronomical research, as it enables discovery of rare or unexpected phenomena within existing observational datasets.
As an interdisciplinary field, ML integrates principles from statistics, optimization theory, and information science. Its primary focus lies in developing adaptive systems that simulate human learning processes to iteratively acquire knowledge, refine skills, and optimize performance. Figure 3 illustrates the iterative workflow of a typical ML system.
@@ 3.1 Supervised Learning
Supervised Learning (SL) operates under guidance, where labeled data provides the necessary supervision for model training. In this framework, class labels act as a reference, enabling the model to learn mappings between inputs and outputs. For instance, in a classification task, such as diagnosing a disease as positive or negative, the algorithm relies on predefined labels to make informed predictions. However, SL is inherently task-specific, meaning the model is limited to learning patterns strictly within the provided data and cannot generalize beyond its training scope Bishop & Nasrabadi (2006). The basic flow of SL is shown in Figure 4.
Figure 4: Basic process of supervised learning.
@@
Decision tree is a non-parametric model built during training, represented as a top-down, tree-like structure. It is used for both classification and regression tasks. The tree consists of sequential nodes, where each node applies a condition to a specific feature in the dataset, guiding the decision-making process Quinlan (1986).

### v2

@@ Abstract
The rapid advancement of observational capabilities in astronomy has led to an exponential growth in the volume of light curve (LC) data, creating both opportunities and challenges for time-domain astronomy. Traditional analytical methods often struggle to fully extract the scientific value of these large and complex datasets. Machine learning (ML) algorithms are increasingly used for LC analysis, enabling classification, prediction, pattern discovery and anomaly detection. However, research in this area remains fragmented, with no comprehensive synthesis of how ML methods address the specific challenges of LC data. Key difficulties include class imbalance, noisy or sparse measurements, effective feature extraction, and limited interpretability of models. This lack of a unified overview makes it difficult for researchers to identify suitable approaches or recognize unresolved problems that require methodological advances. To address this gap, this survey systematically reviews ML techniques applied to LC analysis, outlining their principles and applications in tasks such as exoplanet detection, variable star classification and supernova identification. By clarifying the current state of the field and highlighting open challenges, this work provides guidance for future research and supports the effective integration of ML into astronomical big data studies.
@@ 1 Introduction
Time domain astronomy (TDA) is a rapidly evolving field, driven by the continual discovery of new phenomena and the expansion of observational capabilities Ball & Brunner (2010); Pesenson et al. (2010). This field investigates time-varying characteristics of celestial objects using data from diverse messengers, including gravitational waves, neutrinos, and electromagnetic radiation across various photon-energy bands Vaughan (2013). With the advent of advanced wide-field, multi-epoch sky surveys such as the Sloan Digital Sky Survey (SDSS) York et al. (2000), Zwicky Transient Facility (ZTF) Graham et al. (2019), Panoramic Survey Telescope and Rapid Response System (Pan-STARRS) Kaiser et al. (2010), and the upcoming Vera C. Rubin Observatory (formerly LSST) Ivezić et al. (2019), the volume of transient discoveries has grown exponentially. For instance, ZTF generates alerts for over 100,000 events each night, while LSST is projected to surpass this by an order of magnitude. These alerts, which indicate significant flux density changes or new spatial positions, represent invaluable data streams for studying transient phenomena such as supernovae, pulsars, and gamma-ray bursts. TDA focuses on analyzing time-series data to construct LCs, which depict the variation in brightness of celestial objects over time and are fundamental to understanding the physical processes driving these phenomena.
@@
ML has become an essential tool for automating data classification, archiving, and retrieval Fluke & Jacobs (2020); Baron (2019). Unlike traditional feature extraction methods, which require domain expertise, deep learning (DL) automatically learns data representations. As a subset of ML, DL has gained widespread popularity and has been successfully applied in various domains, including speech recognition, image processing, and natural language processing Nassif et al. (2019); Deng & Li (2013); Rodellar et al. (2018); Mishra & Celebi (2016); Ussipov et al. (2024a, b); Otter et al. (2020). In astronomy, ML has emerged as a powerful tool for automating the classification and analysis of LCs, enabling researchers to extract insights from vast datasets.
Despite its growing adoption, several persistent challenges limit the effectiveness of ML for LC analysis. Class imbalance occurs because some astronomical sources, such as rare transients, are vastly underrepresented compared to more common classes, which can bias models toward majority classes and reduce sensitivity to rare events. Noise arises from observational uncertainties, instrumental limitations, or environmental effects, often degrading model performance and leading to spurious detections. Interpretability is another major challenge, as many ML and DL methods function as “black boxes,” making it difficult to relate predictions to physical processes or diagnose errors. In addition, biases and uncertainties can stem from selection effects in survey design, incomplete or mislabeled training data, and systematic measurement errors; these issues can propagate through the ML pipeline and affect generalizability. Addressing these challenges is crucial for building models that not only produce accurate predictions but also offer trustworthy scientific insights.
@@
This article provides a comprehensive overview of ML algorithms in LC analysis, covering applications such as exoplanet detection, variable star classification, and supernova identification. By presenting a structured examination of ML techniques in this domain, the review aims to foster a deeper understanding of current methodologies and identify chanllenges that should be adressed in the near feature.
@@
The rest of the paper is organized as follows. The methodology is presented in Section 2. Section 3 provides an overview of major photometric data surveys relevant to LC analysis. Section 4 introduces fundamental ML concepts and architectures. Section 5 explores key applications of ML techniques to LC analysis. Section 6 outlines current challenges and open research questions in the field. Finally, Section 7 concludes with a summary and outlook.
2 Review Methodology
To compile this review, we followed the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guidelines Moher et al. (2009). The overall review methodology is illustrated in Figure 1. A comprehensive search was conducted across major scientific databases, including IEEE Xplore, SpringerLink, Elsevier, Web of Science, and Google Scholar, using Boolean combinations of keywords such as “light curve classification”, “light curve machine learning”, “variable star classification”, “supernova classification”, and “exoplanet detection”. The initial search identified 2,780 records, of which 680 duplicates were removed. Title and abstract screening excluded an additional 1,951 records, followed by the removal of 16 records after full-text eligibility assessment. Ultimately, 133 studies were included in this review.
The review covers publications from 2005 to 2025. Studies were included if they (1) were peer-reviewed, (2) applied ML methods to real or simulated LC data, and (3) were relevant to tasks such as variable star classification, exoplanet detection, or transient identification. Non-English papers, duplicate entries, and studies lacking experimental validation or methodological clarity were excluded.
Figure 1: PRISMA flowchart of review methodology.
@@ 3 Photometric Survey Datasets
Over the past few decades, large-scale photometric surveys have transformed time-domain astronomy, providing vast, multi-wavelength datasets that are readily accessible through online databases. These surveys have been essential in detecting and classifying variable stars, discovering exoplanets, and identifying transient astronomical events. As shown in Figure 2, the data volume from astronomical surveys has grown exponentially, approximately doubling every 16 months. The combination of different survey datasets enables comprehensive studies of stellar variability across various timescales and wavelengths, driving both traditional astrophysical research and modern ML applications. This rapid data growth has created a strong demand for automated and scalable analysis methods, especially those based on ML.
@@
Understanding the properties of these surveys is essential for ML applications, since survey-specific factors such as cadence, photometric precision, time coverage, and instrumental effects directly affect the structure and quality of LC data. In addition, biases and uncertainties introduced during survey design or data processing can influence ML model performance, generalization, and reliability. Summarizing major photometric surveys provides important context for selecting appropriate ML techniques and for understanding the practical limitations of real-world astronomical data.
Photometric surveys can be broadly categorized into ground-based and space-based missions. Ground-based surveys provide wide-field and high-cadence monitoring, often covering extensive regions of the sky over long time periods. In contrast, space-based missions offer high-precision, uninterrupted observations unaffected by atmospheric distortions, enabling the detection of minute variations in stellar brightness. Table LABEL:table1 provides an overview of major ground-based and space-based photometric surveys, including their active years, observed bands, sky coverage, and references.
@@ 4 Machine learning fundamentals
Figure 3: Overview of main ML approaches.
Figure 4: Basic process of ML. Cylinder represents database, rectangles indicate data processing steps, and the rectangle with a folded corner denotes knowledge extraction.
ML enables systems to autonomously learn patterns from data and improve decision-making through experience, mirroring aspects of human cognition. Unlike traditional astronomical programming, which relies on explicit physical rules, ML algorithms derive implicit relationships directly from observational data. This data-driven approach offers flexibility for solving complex, nonlinear problems that defy conventional analytical methods Rodr´ıguez et al. (2022); Kembhavi & Pattnaik (2022); Sen et al. (2022).
ML algorithms can be broadly categorized into supervised and unsupervised methods, often referred to as predictive and descriptive, respectively. These approaches may also be combined to form semi-supervised methods. Figure 3 provides a visual taxonomy of these approaches. Supervised algorithms learn mappings between input features and predefined target variables using labeled training data curated by domain experts Zucker & Giryes (2018); Delli Veneri et al. (2019); Krone-Martins et al. (2018); Mahabal et al. (2019); D’Isanto et al. (2018); Norris et al. (2019). Unsupervised methods, conversely, autonomously identify hidden structures or relationships within unlabeled datasets. These techniques are typically divided into three subcategories: clustering (grouping similar data points), dimensionality reduction (extracting salient features), and anomaly detection (identifying outliers) Hocking et al. (2015); Nun et al. (2016); Gianniotis et al. (2016); Baron & Poznanski (2017); Reis et al. (2018a, b). Anomaly detection holds particular promise for astronomical research, as it enables discovery of rare or unexpected phenomena within existing observational datasets.
Another major category is reinforcement learning, which is based on learning through interaction with an environment to maximize cumulative rewards. However, it is excluded from the scope of this review, as its application to LC analysis remains limited due to the lack of interactive, feedback-driven settings in typical observational astronomy workflows.
As an interdisciplinary field, ML integrates principles from statistics, optimization theory, and information science. Its primary focus lies in developing adaptive systems that simulate human learning processes to iteratively acquire knowledge, refine skills, and optimize performance. Figure 4 illustrates the iterative workflow of a typical ML system. Knowledge extraction from data can be effectively achieved using established methodologies such as the Knowledge Discovery in Databases (KDD) process Fayyad et al. (1996) and the Cross-Industry Standard Process for Data Mining (CRISP-DM) Wirth & Hipp (2000).
@@ 4.1 Supervised Learning
Supervised learning (SL) operates under guidance, where labeled data provides the necessary supervision for model training. In this framework, class labels or continuous target values act as references, enabling the model to learn mappings between inputs and outputs. The basic flow of SL is shown in Figure 5. SL problems are typically divided into classification and regression. Classification involves predicting discrete class labels (e.g., variable stars based on their LC), while regression involves predicting continuous values (e.g., estimating a star’s period or temperature from its LC). Both approaches have been applied to LC analysis: classification is widely used for star type identification, supernova classification and exoplanet detection Akhmetali et al. (2024); Mahabal et al. (2017); Valizadegan et al. (2022); Pearson et al. (2018); Mitra et al. (2023, 2024); Abylkairov et al. (2025); Bissekenov et al. (2024), whereas regression is used for parameter estimation and LC detrending Hinners et al. (2018); Thieler et al. (2016); Aigrain et al. (2016); Paunzen & Vanmunster (2016); Carter & Winn (2009).
Figure 5: Basic process of SL. Cylinder represents database, rectangles indicate data processing steps, parallelograms denote datasets, and the rounded rectangle shows the final output.